OpenStudy (anonymous):

Solve by factoring the trinomial as a binomial square and then use the square root method. Round to the nearest tenth if necessary. x2 + 16x + 64 = 81 A. {9, -9} B. {-8, 9} C. {-17, 1} D. no real solution

OpenStudy (anonymous):

\[x^2+16x+64=81\] \[(x+4)^2=81\] \[x+4=9\] or \[x+4=-9\] \[x=5\] or \[x=-13\]

OpenStudy (anonymous):

...... wait.... ?

OpenStudy (anonymous):

i am wrong sorry

OpenStudy (anonymous):

\[(x+8)^2=81\] \[x+8=9\] \[x+8=-9\] \[x=1\] \[x=-17\]

OpenStudy (anonymous):

that is what i get for rushing.

OpenStudy (anonymous):

Haha (: Thanks. Your so effing helpful!!!!!!!!

OpenStudy (anonymous):

must be getting tired.

OpenStudy (anonymous):

promise i will be more careful

OpenStudy (anonymous):

=3 Ahww! I am totally getting free answers from you!! And your totally being nice!! =3 If i didnt have a boyfriend i would stalk you .. lmfao totally kidding ^^ i just like to use my name to my advantage ^^

OpenStudy (anonymous):

O-o_Be-Creepy_o_O i cannot even type it. where did it come from?

OpenStudy (anonymous):

Haha idk i used my blonde imagination ^_^ im very creative!! (:

hero (hero):

Wow, this is old

hero (hero):

looks like somebody misses creepy :D